MECHANISM OF THE MULTIPHOTON DISSOCIATION OF SO_{2} VIA THE H̃ RYDBERG STATE

\underline{YUXIU LEI}, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104; ED-WIN QUIÑONES, Department of Chemistry, University of Puerto Rico, San Juan, PR 00931; AMI-TAVIKRAM A. DIXIT, KEON WOO LEE, PAUL L. HOUSTON, Department of Chemistry and Chemical Biology, Cornell University, Ithaca, NY 14853.

We will present a study on the mechanism of multiphoton dissociation of the SO_{2} molecule through the H̃ Rydberg state. For this study, a homemade Wiley-McLaren linear time-of-flight (TOF) spectrometer was constructed to characterize the fragments produced during irradiation of SO_{2} with a laser beam in the 222 - 234 nm region. Both SO^{+} and S^{+} fragment ions were detected but no parent SO_{2}^{+} appeared. The mass resolved excitation spectra (MRES) of both SO^{+} and S^{+} display the SO_{2} C̃ ^{1}B_{2} ← X^{1}A_{1} vibrational progressions. The first photon in the studied region pumps SO_{2} to the C̃ ^{1}B_{2} electronic state. The second photon further excites the molecule to the H̃ Rydberg state, where it dissociates into SO + O and S + O_{2} fragments. The third photon then ionizes the SO fragment. From the power dependence of the ion intensity and a kinetic model we developed, we conclude that S^{+} is produced by a one-photon dissociation, SO^{+} + hν → S^{+} + O. The internal energy distributions of the fragmentation products from dissociation of SO_{2} via the H̃ Rydberg state were also obtained using an ion-imaging spectrometer. The following five dissociation channels are observed: SO(B ^{3}Σ^{−}) + O(^{3}P_{2}), SO(A^{3}Π) + O(^{3}P_{2}) , S(^{1}D) + O_{2}(b^{1}Σ_{g}), S(^{1}D) + O_{2}(a^{1}Δ_{g}), and S(^{1}D) + O_{2}(X^{3}Σ^{−}).